# RLHF Workflow: From Reward Modeling to Online RLHF

### v2

@@ 3.3 Practical Implementation Details
To facilitate exploration, we combine the temperature tuning with the rejection sampling strategy with $n=8$. Instead of fixing $\pi_{t}^{1}=\pi^{\mathrm{MLE}}_{t}$ (like the center of confidence set) and optimizing the $\pi_{t}^{2}$ solely to be the best-of-8 variant of $\pi^{\mathrm{MLE}}_{t}$, we take $\pi_{t}^{1}$ and $\pi_{t}^{2}$ as the best-of-8 policy and worst-of-8 policy induced by $\pi^{\mathrm{MLE}}_{t}$. In other words, we take the best response and the worst response as ranked by the reward model to get a preference pair. In this case, we jointly optimize the two policies to maximize their difference (measured by the uncertainty), which tends to be more efficient in practice and enjoys the same theoretical guarantee as stated in Theorem 1. This choice is similar to Hoang Tran, (2024); Pace et al., (2024); Yuan et al., 2024b ; Xu et al., (2024). We also drop the pair where $\pi_{t}^{1}$ and $\pi_{t}^{2}$ give the same response, which implies that the uncertainty in this direction is already small. For this round of experiments, we still use the reward function trained as the MLE of the BT reward model to rank the responses for the following reasons. First, to rank $n$ responses, the complexity of using the reward model is linear in $n$, while it is far more complicated with the pairwise preference model. Second, during the early experiments, we observe significant length bias in the iterative RLHF. Therefore, we would like to explore the strategy to mitigate the length bias, and it is relatively easier to penalize the reward value with the length of the response. Finally, the BT reward model is comparable with the preference model except for the reasoning task and it may be already satisfactory for our goal. We leave a more comprehensive comparison between the BT reward model and preference model for future study.
@@
Prompt set, and data generation. We collect prompts from UltraFeedback (Cui et al.,, 2023), HelpSteer (Wang et al.,, 2023), OpenOrca (Lian et al., 2023a, ), UltraInteract (Yuan et al., 2024a, ), Capybara (Daniele and Suphavadeeprasit,, 2023) and DIBT-10K [^6] https://huggingface.co/datasets/DIBT/10k_prompts_ranked and prepare the full prompt set here [^7] https://huggingface.co/datasets/RLHFlow/prompt-collection-v0.1. In our experiments, we use a subset of 60K prompts and iterate for three iterations, so 20K prompts are used to generate 20K x 8 responses per iteration. To accelerate data generation, we use VLLM (Kwon et al.,, 2023) for inference. We set the max generation length as 2048, and use a sampling temperature of 1.0/0.7 without any top-k/top-p strategy. To offer a more intuitive comprehension of our prompts collection, we provide visualization plots in Figure 7, which are generated on Nomic Atlas [^8] https://atlas.nomic.ai/ with the nomic-embed-text-v1.5 text embedding model (Nussbaum et al.,, 2024).
@@ 4 Evaluation of the Model
4.1 Benchmark
We evaluate the models by standard benchmarks, including AlpacaEval-2, MT-Bench, and Chat-Arena-Hard.
  • AlpacaEval-2 (Dubois et al.,, 2023): This benchmark focuses on single-turn conversations and consists of 805 test prompts covering various topics. The models are compared head-to-head with GPT-4-Preview (11/06) to compute the win rate. The same GPT-4 model is used as the judge. To mitigate the length bias of GPT-4, a length-control variant of the benchmark is also proposed.
  • MT-Bench (Zheng et al.,, 2023): This benchmark is a multi-turn benchmark and includes 160 test prompts from 8 different areas. The model should first answer an initial question, and then a pre-defined follow-up question. The model’s responses are then rated by the GPT-4 model with a scale from 1-10, and the final score is computed as the average score of two turns.
  • Chat-Arena-Hard (Tianle et al.,, 2024): This benchmark consists of 500 test prompts from the live data in Chatbot Arena, a crowd-sourced platform for LLM evaluations. The prompts evaluate the model’s ability in specificity, domain knowledge, complexity, problem-solving, creativity, technical accuracy, and real-world application. In addition to the agreement to human preference, compared with AlpacaEval-2 and MT-Bench, Chat-Arena-Hard further enjoys a clear separability among different models.
We also measure the ability of the resulting models using academic benchmark, including GSM-8K (Cobbe et al.,, 2021), MMLU (Hendrycks et al.,, 2020), HumanEval (Chen et al.,, 2021), TruthfulQA (Lin et al.,, 2021), ARC (Clark et al.,, 2018), and MBPP (Austin et al.,, 2021). These benchmarks evaluate the models’ ability in coding, reasoning, and general knowledge. In particular, it is known that RLHF alignment can introduce performance degeneration in reasoning, calibration (providing accurate confidence estimates), and truthfulness capabilities (generating accurate and factual responses), which is also referred to as the alignment tax in the literature (Ouyang et al.,, 2022; Bai et al., 2022a, ; OpenAI,, 2023). Therefore, evaluating our model on these benchmarks is crucial to understanding the impact of iterative RLHF on these specific aspects.
We summarize the benchmarks we use in this project in Table 3.
Metric win rate score win rate acc acc acc acc acc acc
Num. of Shots 0 0 0 8 5 0 0 25 0
@@ 4.2 Main Results
Table 4: Evaluation results and comparison between the resulting models and existing models. means that the model is based on the mixture-of-experts architecture. We report the length-control win rate of AlpacaEval-2 as recommended by the authors. RS is short for rejection sampling (Dong et al.,, 2023) and X means that the value is unavailable. Only underline results are better than our 8B model.
@@
Online iterative RLHF significantly improves conversation quality. We evaluate our model’s conversation abilities using AlpacaEval-2, MT-Bench, and Chat-Arena-Hard, (results in Table 4). Compared to other open-source models with less than 10B parameters, our model – LLaMA-3-8B-SFR-Iterative-DPO-R outperforms them on the conversation and instruction-following benchmarks with a significant margin. Notably, our model trained with iterative DPO consistently outperforms that of vanilla offline DPO (DPO baseline). This demonstrates the advantage of online iterative RLHF. Moreover, our model outperforms the Tulu-2-DPO-70B and GPT-3.5-turbo-1106, which are aligned by DPO or PPO and are much larger than our base model. These results show that the online iterative RLHF can effectively adjust the style of the model responses, thus improving the conversation quality.
Academic Task. As RLHF can impact a model’s reasoning and calibration abilities, typically in a negative way (Bai et al., 2022a, ; Ouyang et al.,, 2022; OpenAI,, 2023), we compare our model’s performance on academic benchmarks (Table 5) with the SFT checkpoint and other baselines. We don’t observe significant performance regression compared to the SFT baseline. Interestingly, our iteratively DPO-aligned model even outperforms the SFT model in GSM-8K, MMLU, TruthfulQA, and ARC benchmarks. We believe that these increased capacities of the model are injected in the pre-training stage and SFT stage, and iterative DPO helps it leverage them more effectively. This is because the 60K alignment data used in the iterative RLHF are orders of magnitude less than those used in the previous two stages.
@@ 5 End Note and Future Direction
In this technical report, we study the workflow of the online iterative RLHF, which leverages on-policy sampling and external preference signals from a proxy preference model trained on a diverse set of open-source preference datasets. The resulting model (referred to as LLaMA-3-8B-SFR-Iterative-DPO-R) demonstrates impressive performance on standard benchmarks, and the report provides detailed instructions for reproducing the results, including data, code, models, and hyper-parameter choices.
There are still many potential directions to explore. First, as we can see in Table 6, the iterative RLHF heavily relies on the quality of the preference signal. In this project, we use a proxy scalar reward model trained on a diverse set of open-source datasets to approximate human feedback. It would be interesting to see whether we can design a more effective strategy to model different types of preference signals, like a multi-head reward (Wang et al.,, 2024) and classification-based activation strategy (Touvron et al.,, 2023). Second, while the rejection sampling seems to be a good heuristic exploration strategy, it is still interesting to see whether we can design more effective ways for exploration. Finally, most of the models after RLHF tend to reply the prompts with much longer responses. Such a length bias is further amplified in the iterative RLHF framework. We presented a preliminary study on this issue by leveraging an additional length penalty in reward for data filtering. It would be interesting to see whether we can further mitigate this issue by additional algorithmic designs or post-training techniques.
@@ Appendix B Additional Experimental Details
SFT Data List. We collect open-sourced instruction-finetuning data for our SFT model training. The following data is included: ShareGPT, Evol-Instruct, SlimOrca, MathInstruct, Magicoder-Evol-Instruct, GPT4-LLM, OrcaMath, GPTeacher, UltraInteract.
@@
Offline Vanilla DPO. We use Nectar dataset for Offline DPO. We run 1 epoch with batch size 128, learning rate 5e-7, and cosine decay scheduler.
Evaluation. For MT-Bench evaluation, we use huggingface default generation and “You are CMB, an AI assistant known for its intelligence and expertise across all fields of knowledge. You are designed to provide detailed and helpful responses to a wide range of user inquiries, ensuring clarity and accuracy in every interaction.” as system prompt. Transformers version is 4.38.1. For AlpacaEval and Chat-Arena-Hard, we use vllm to generate samples.
@@
Additional Plots. We also visualized our performance as Figure 8.

### v3

@@ 3.3 Practical Implementation Details
To facilitate exploration, we combine the temperature tuning with the rejection sampling strategy with $n=8$. Instead of fixing $\pi_{t}^{1}=\pi^{\mathrm{MLE}}_{t}$ (like the center of confidence set) and optimizing the $\pi_{t}^{2}$ solely to be the best-of-8 variant of $\pi^{\mathrm{MLE}}_{t}$, we take $\pi_{t}^{1}$ and $\pi_{t}^{2}$ as the best-of-8 policy and worst-of-8 policy induced by $\pi^{\mathrm{MLE}}_{t}$. In other words, we take the best response and the worst response as ranked by the reward model to get a preference pair. In this case, we jointly optimize the two policies to maximize their difference (measured by the uncertainty), which tends to be more efficient in practice and enjoys the same theoretical guarantee as stated in Theorem 1. This choice is similar to Hoang Tran (2024); Pace et al. (2024); Yuan et al. (2024b); Xu et al. (2024). We also drop the pair where $\pi_{t}^{1}$ and $\pi_{t}^{2}$ give the same response, which implies that the uncertainty in this direction is already small. For this round of experiments, we still use the reward function trained as the MLE of the BT reward model to rank the responses for the following reasons. First, to rank $n$ responses, the complexity of using the reward model is linear in $n$, while it is far more complicated with the pairwise preference model. Second, during the early experiments, we observe significant length bias in the iterative RLHF. Therefore, we would like to explore the strategy to mitigate the length bias, and it is relatively easier to penalize the reward value with the length of the response. Finally, the BT reward model is comparable with the preference model except for the reasoning task and it may be already satisfactory for our goal. We leave a more comprehensive comparison between the BT reward model and preference model for future study.
@@
Prompt set, and data generation. We collect prompts from UltraFeedback (Cui et al., 2023), HelpSteer (Wang et al., 2023), OpenOrca (Lian et al., 2023a), UltraInteract (Yuan et al., 2024a), Capybara (Daniele & Suphavadeeprasit, 2023) and DIBT-10K [^4] https://huggingface.co/datasets/DIBT/10k_prompts_ranked and prepare the full prompt set.In our experiments, we use a subset of 60K prompts and iterate for three iterations, so 20K prompts are used to generate 20K x 16 responses per iteration. To accelerate data generation, we use VLLM (Kwon et al., 2023) for inference. We set the max generation length as 2048, and use a sampling temperature of 1.0/0.7 without any top-k/top-p strategy. To offer a more intuitive comprehension of our prompts collection, we provide visualization plots in the Appendix (Figure 5).
@@ 4 Evaluation of the Model
4.1 Benchmarks
We evaluate the models by standard benchmarks, including AlpacaEval-2, MT-Bench, and Chat-Arena-Hard. Details are provided in the Appendix.
We also measure the ability of the resulting models using academic benchmark, including GSM-8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2020), HumanEval (Chen et al., 2021), TruthfulQA (Lin et al., 2021), ARC (Clark et al., 2018), and MBPP (Austin et al., 2021). These benchmarks evaluate the models’ ability in coding, reasoning, and general knowledge. In particular, it is known that RLHF alignment can introduce performance degeneration in reasoning, calibration (providing accurate confidence estimates), and truthfulness capabilities (generating accurate and factual responses), which is also referred to as the alignment tax in the literature (Ouyang et al., 2022; Bai et al., 2022a; OpenAI, 2023). Therefore, evaluating our model on these benchmarks is crucial to understanding the impact of iterative RLHF on these specific aspects.
@@ 4.2 Main Results
Table 2: Evaluation results and comparison between the resulting models and existing models. means that the model is based on the mixture-of-experts architecture. We report the length-control win rate of AlpacaEval-2 as recommended by the authors. RS is short for rejection sampling (Dong et al., 2023) and X means that the value is unavailable. Only underline results are better than our 8B model.
@@
Online iterative RLHF significantly improves conversation quality. We evaluate our model’s conversation abilities using AlpacaEval-2, MT-Bench, and Chat-Arena-Hard, (results in Table 2). Compared to other open-source models with less than 10B parameters, our model outperforms them on the conversation and instruction-following benchmarks with a significant margin. Notably, our model trained with iterative DPO consistently outperforms that of vanilla offline DPO (DPO baseline). This demonstrates the advantage of online iterative RLHF. Moreover, our model outperforms the Tulu-2-DPO-70B and GPT-3.5-turbo-1106, which are aligned by DPO or PPO and are much larger than our base model. These results show that the online iterative RLHF can effectively adjust the style of the model responses, thus improving the conversation quality.
Academic Task. As RLHF can impact a model’s reasoning and calibration abilities, typically in a negative way (Bai et al., 2022a; Ouyang et al., 2022; OpenAI, 2023), we compare our model’s performance on academic benchmarks (Table 3) with the SFT checkpoint and other baselines. We don’t observe significant performance regression compared to the SFT baseline. Interestingly, our iteratively DPO-aligned model even outperforms the SFT model in GSM-8K, MMLU, TruthfulQA, and ARC benchmarks. We believe that these increased capacities of the model are injected in the pre-training stage and SFT stage, and iterative DPO helps it leverage them more effectively. This is because the 60K alignment data used in the iterative RLHF are orders of magnitude less than those used in the previous two stages.
@@ 5 End Note and Future Direction
In this technical report, we study the workflow of the online iterative RLHF, which leverages on-policy sampling and external preference signals from a proxy preference model trained on a diverse set of open-source preference datasets. The resulting model demonstrates impressive performance on standard benchmarks, and the report provides detailed instructions for reproducing the results, including data, code, models, and hyper-parameter choices.
There are still many potential directions to explore. First, as we can see in Table 4, the iterative RLHF heavily relies on the quality of the preference signal. In this project, we use a proxy scalar reward model trained on a diverse set of open-source datasets to approximate human feedback. It would be interesting to see whether we can design a more effective strategy to model different types of preference signals, like a multi-head reward (Wang et al., 2024) and classification-based activation strategy (Touvron et al., 2023). Second, while the rejection sampling seems to be a good heuristic exploration strategy, it is still interesting to see whether we can design more effective ways for exploration. Finally, most of the models after RLHF tend to reply the prompts with much longer responses. Such a length bias is further amplified in the iterative RLHF framework. We presented a preliminary study on this issue by leveraging an additional length penalty in reward for data filtering. It would be interesting to see whether we can further mitigate this issue by additional algorithmic designs or post-training techniques.
@@
Update (Nov. 2024). We also provide an updated version of SFT dataset [^5] https://huggingface.co/datasets/RLHFlow/RLHFlow-SFT-Dataset-ver2 and the corresponding model [^6] https://huggingface.co/RLHFlow/LLaMA3-SFT-v2.
@@ Appendix B Additional Experimental Details
B.1 Preference Datasets
Following the Pair-RM (Jiang et al., 2023) and LLaMA-2 (Touvron et al., 2023), we use a mixture of open-source datasets as the training set. Here is a brief introduction to the datasets:
  • HH-RLHF (Bai et al., 2022a) is a pairwise preference dataset where each sample is accompanied by a conversation history and two alternative responses written by an early Claude model with 52B parameters. The preferences of the responses are annotated by humans.
  • SHP (Ethayarajh et al., 2022) is sourced from Reddit and includes examples from 18 subreddits, such as askacademia, askbaking, askengineers, and changemyview. Each example is a Reddit post with a question/instruction and a pair of top-level comments. One comment is preferred by more Reddit users than the other. All preferences and responses are provided by humans. Following Cui et al. (2023), only samples with a score ratio $>$ 2 are used, and at most 5 pairs are taken for each prompt.
  • HelpSteer (Wang et al., 2023). This open-source dataset (Wang et al., 2023) contains prompts, responses, and five human-annotated attributes (helpfulness, correctness, coherence, complexity, and verbosity) ranging from 0 to 4. The prompts are generated using a mixture of template-generated and human-generated methods, while responses are generated by an in-house LLM. The authors generate up to 4 responses per prompt, and we can construct pairwise comparisons based on them.
  • PKU-SafeRLHF (Ji et al., 2024). This dataset (Ji et al., 2024) consists of 30k+ expert comparison data. Each sample includes two responses to a question and two preference signals for helpfulness and safety, respectively. The responses are generated by open-source chatbots, and the preference signals are merged through the results of 14 harm category multi-class classficiation.
  • UltraFeedback (Cui et al., 2023) consists of 64k prompts from diverse resources (including UltraChat, ShareGPT, Evol-Instruct, TruthfulQA, FalseQA, and FLAN) and the authors generate 4 responses per prompt using 4 different LLMs sampled from a diverse set of state-of-the-art open-source LLMs. The preference is from GPT-4 based on a fine-grained annotation instruction, which contains 4 different aspects, namely instruction-following, truthfulness, honesty and helpfulness. The dataset collection strategy of UltraFeedback has also influenced many subsequent works.
  • CodeUltraFeedback is generated similarly with the Ultrafeedback but focuses on the coding task. The annotation is from GPT-3.5.
  • UltraInteract (Yuan et al., 2024a) is a preference dataset designed for complex reasoning tasks. The authors collect a preference tree for each instruction, with the instruction being the root and each action a node. A trajectory is a root-to-leaf path consisting of a sequence of actions. Paired correct and incorrect nodes or trajectories are used for preference learning.
  • Distilabel-Capybara [^7] https://huggingface.co/datasets/argilla/distilabel-capybara-dpo-7k-binarized is a preference dataset of multi-turn dialogues whose prompts are taken from Daniele & Suphavadeeprasit (2023), where the responses are generated by open-source LLMs and preferences are generated by GPT-4.
  • Distilabel-Orca [^8] https://huggingface.co/datasets/argilla/distilabel-intel-orca-dpo-pairs is collected similarly with Capybara but with the prompts from Lian et al. (2023a).
HH-RLHF 115092 160.4 82.2 73.6 LLM Human 115396
SHP 31003 186.2 173.6 88.8 Human Human 93301
HelpSteer 8592 530 116.4 89.3 LLM Human 37131
PKU-SafeRLHF-30K 6975 21.5 70.4 74.6 LLM Human 26874
UltraFeedback 63591 161.5 279.5 211.1 LLM GPT-4 340025
UltraInteract 76086 507.4 396.6 416.7 LLM GPT-4 161927
CodeUltraFeedback 9938 172.8 427.6 400.6 LLM GPT-3.5 50156
Argilla-Math 2352 36.5 276.5 265.3 LLM GPT-4 2418
OpenOrca 6791 153.3 165.4 260.5 LLM GPT-4 6926
Capybara 14740 634.5 348.4 401.9 LLM GPT-4 14811
The training of LLMs is highly data-dependent. To ensure high-quality training data, we conduct a filtering process on the open-source datasets we use. This process removes low-quality and meaningless samples. Additionally, conversations with empty rounds or incorrect labels (implied by the other features of the dataset) are eliminated. Furthermore, in datasets where absolute scores are available, pairwise comparisons with small margins are excluded as these preference signals tend to be noisy (Bansal et al., 2023). This process roughly deletes 10% of the data. We summarize the statistics of the open-source datasets that are used for the training in Table 5 and prepare them, as well as our data filtering script, on the huggingface.
We consider two versions of the training set:
  • Mix1: HH-RLHF + SHP + UltraFeedback + Summarization (Stiennon et al., 2020).
  • Mix2: all the datasets in Table 5.
The Mix1 dataset is similar to the construction of UltraFeedback (Cui et al., 2023) with an additional summarization dataset. In comparison, the Mix2 consists of more reasoning preference pairs (math and code) and safety data. We also consider three different approaches to model the preference signals, including prompting in the LLM-as-a-judge manner (Zheng et al., 2023), reward modeling as the MLE of the BT reward model, and the preference model.
B.2 Benchmark Details
  • AlpacaEval-2 (Dubois et al., 2023): This benchmark focuses on single-turn conversations and consists of 805 test prompts covering various topics. The models are compared head-to-head with GPT-4-Preview (11/06) to compute the win rate. The same GPT-4 model is used as the judge. To mitigate the length bias of GPT-4, a length-control variant of the benchmark is also proposed.
  • MT-Bench (Zheng et al., 2023): This benchmark is a multi-turn benchmark and includes 160 test prompts from 8 different areas. The model should first answer an initial question, and then a pre-defined follow-up question. The model’s responses are then rated by the GPT-4 model with a scale from 1-10, and the final score is computed as the average score of two turns.
  • Chat-Arena-Hard (Tianle et al., 2024): This benchmark consists of 500 test prompts from the live data in Chatbot Arena, a crowd-sourced platform for LLM evaluations. The prompts evaluate the model’s ability in specificity, domain knowledge, complexity, problem-solving, creativity, technical accuracy, and real-world application. In addition to the agreement to human preference, compared with AlpacaEval-2 and MT-Bench, Chat-Arena-Hard further enjoys a clear separability among different models.
We also summarize the benchmarks we use in this project in Table 6.
Metric win rate score win rate acc acc acc acc acc acc
Num. of Shots 0 0 0 8 5 0 0 25 0
B.3 Other details
Prompt Visualization.
We provide the visualization generated on Nomic Atlas [^9] https://atlas.nomic.ai/ with the nomic-embed-text-v1.5 text embedding model (Nussbaum et al., 2024)
SFT Data List.
@@ B.3 Other details
We collect open-sourced instruction-finetuning data for our SFT model training. The following data is included: ShareGPT (Chiang et al., 2023), Evol-Instruct (Xu et al., 2023a), SlimOrca (Lian et al., 2023b), MathInstruct (Yue et al., 2023), Magicoder-Evol-Instruct (Wei et al., 2023), GPT4-LLM (Peng et al., 2023), OrcaMath (Mitra et al., 2024), GPTeacher (Teknium1, 2023), UltraInteract (Yuan et al., 2024a).
Offline Vanilla DPO.
@@
We use Nectar dataset for Offline DPO. We run 1 epoch with batch size 128, learning rate 5e-7, and cosine decay scheduler.
Additional Plots.
@@
We also have some additional training plots Figure 6, 7 and have visualized our performance as Figure 9.
Hyperparameters.
The hyperparameters are listed in Table 7.
Parameter Value
n_batch_size_per_device 2
n_gradient_accumulation 8
optim adamw_torch
lr_scheduler_type cosine
num_train_epochs 2
beta 0.1
Table 7: Training parameters
Figure 8: Model performance with respect to RLHF iterations.
@@
Appendix C Case Studies
To showcase the significant improvements achieved through online RLHF, we conducted a qualitative analysis of the generated responses. We observed that, after applying online RLHF, the responses became more detailed and better formatted, often utilizing bullet points, highlights, bold text, and enumeration. This enhanced clarity and structure were key factors contributing to the improved win-rate observed in our experiments. Below, we present several examples that illustrate these improvements.
Prompt What is Gremolata?
Offline RLHF Response Gremolata is a flavorful, aromatic mixture used as a garnish or seasoning for a variety of Italian dishes, particularly osso buco, a dish made from braised veal shanks. It is also used in other dishes like fish, pasta, or vegetables. Gremolata is typically made from a combination of finely chopped fresh parsley, garlic, and lemon zest. Sometimes, basil or mint can be added as well. The mixture is usually prepared by finely chopping or mincing the ingredients and then mixing them together. Gremolata adds a bright, fresh, and citrusy flavor to the dishes it is used in. It is often sprinkled on top of the dish just before serving, allowing the flavors to meld and enhance the overall taste.
Online RLHF Response Gremolata is a flavorful, aromatic mixture used as a garnish or seasoning in Italian cuisine, particularly in dishes such as Osso Buco, risotto, and other meat or fish dishes. It is often served on top of the dish to add a burst of fresh flavors. The basic ingredients of Gremolata are: 1. Fresh parsley: This is the main component, providing the fresh, herbaceous flavor and color to the mixture. Flat-leaf parsley is often preferred for its robust flavor. 2. Garlic: Finely chopped or minced raw garlic gives a sharp, pungent taste to the Gremolata. 3. Lemon zest: The zest from a lemon adds a bright, citrusy flavor and a slight tang to the mixture. In some variations, you might also find: 4. Lemon juice: This adds a little extra acidity and freshness to the mix. 5. Fresh basil or mint: These herbs can be used instead of, or in addition to, parsley, depending on the dish and personal preference. To prepare Gremolata, simply mix the ingredients together in a small bowl. The proportions can vary depending on personal taste, but a general guideline is 1 part lemon zest, 2 parts chopped parsley, and 12 part chopped garlic. Adjust to your liking. Gremolata is typically added just before serving to maintain its freshness and bright flavors. It pairs well with rich and savory dishes, as it helps to brighten and lighten the overall taste.
Figure 10: Case Study – Example 1.
Prompt What are the best exercises for beginners?
Offline RLHF Response The best exercises for beginners are those that are simple, safe, and effective in improving overall fitness. Some of the top beginner exercises include: 1. Squats: Helps to strengthen the lower body, particularly the glutes, quadriceps, and hamstrings. 2. Push-ups: Improves upper body strength, targeting the chest, triceps, and shoulders. 3. Planks: Engages the core muscles and strengthens the lower back, arms, and glutes. 4. Lunges: Strengthens the quadriceps, hamstrings, and glutes, while improving balance and flexibility. 5. Standing Rows: A great exercise for the upper back, biceps, and shoulders. 6. Bicep curls: Targets the biceps, promoting arm strength and definition. 7. Tricep extensions: Strengthens the triceps, an important muscle for overall upper body strength. 8. Walking or jogging: A simple and effective way to improve cardiovascular fitness and burn calories. 9. Bodyweight squats and lunges: These are great for beginners as they only require bodyweight and help build strength and balance. 10. Stretching and yoga: Incorporating flexibility exercises like yoga or simple stretches can help prevent injury and promote overall mobility. Remember to start with low weights or bodyweight exercises, gradually increasing the intensity, and always focus on proper form to avoid injury. It’s also a good idea to start with a warm-up and cool-down, and consider working with a personal trainer or coach to learn the correct techniques.
Online RLHF Response The best exercises for beginners are those that are simple, safe, and effective in improving overall fitness, strength, and flexibility. Here are some top recommendations for starting your fitness journey: 1. **Bodyweight Squats**: A fundamental exercise that targets the lower body, including the glutes, hamstrings, and quadriceps. Stand with feet shoulder-width apart, lower your hips as if sitting back into a chair, then push back up to the starting position. Aim for 2-3 sets of 10-15 reps. 2. **Push-ups**: A classic upper body strengthener that targets the chest, triceps, and shoulders. Start on your hands and knees, then lower your body until your chest nearly touches the ground, then push back up. Modify by starting on your knees if needed. Aim for 2-3 sets of 8-12 reps. 3. **Plank**: A core-strengthening exercise that also works the shoulders, back, and glutes. Start in a push-up position, then lower your forearms to the ground, keeping your body in a straight line from head to heels. Hold for 30 seconds to 1 minute, then rest and repeat for 2-3 sets. 4. **Lunges**: A lower body exercise that targets the glutes, hamstrings, and quadriceps. Stand with feet together, then step forward with one foot, lowering your hips until both knees are bent at a 90-degree angle. Push back to the starting position and repeat on the other side. Aim for 2-3 sets of 8-12 reps per leg. 5. **Bent-over Row**: A back and shoulder strengthener that targets the latissimus dorsi, rhomboids, and traps. Hold a dumbbell in each hand, bend forward at the hips with a straight back, then row the weights up to your sides, squeezing your shoulder blades together. Lower the weights and repeat for 2-3 sets of 8-12 reps. 6. **Walking or Jogging**: A great way to improve cardiovascular fitness and burn calories. Start with a brisk walk, gradually increasing speed and duration as you become more comfortable. Aim for at least 30 minutes of moderate-intensity exercise, such as brisk walking, 5 days a week. 7. **Yoga or Pilates**: These low-impact exercises are excellent for improving flexibility, balance, and core strength. Start with beginner classes or videos, focusing on proper form and breathing. Aim for at least 2-3 sessions per week. 8. **Calf Raises**: A lower body exercise that targets the calf muscles. Stand on the edge of a step or platform, then raise your heels as high as possible, then lower back down. Aim for 2-3 sets of 10-15 reps. Remember to warm up before starting any exercise, and always listen to your body. If an exercise causes pain or discomfort, stop and consult a healthcare professional or fitness instructor. As you progress, you can gradually increase the intensity, duration, and variety of your workouts to continue challenging your body and achieving your fitness goals.
Figure 11: Case Study – Example 2.
Prompt "What type of soil is suitable for cactus?"
Offline RLHF Response Cactus plants thrive in well-draining soil that is porous and allows excess water to escape quickly. The ideal soil for cactus should have a pH level between 6.0 and 7.5, which is slightly acidic to neutral. The soil should also be low in nutrients, as cactus plants generally prefer a nutrient-poor environment. The following soil types are suitable for cactus plants: 1. Cactus-specific soil mix: This is a commercial soil mix designed specifically for cactus plants. It typically contains a blend of sand, coconut coir, perlite, and peat moss or sphagnum moss. This mix provides excellent drainage and aeration while retaining moisture. 2. Desert soil mix: A desert soil mix can also be suitable for cactus plants. It usually consists of a blend of sand, perlite, andor vermiculite. This mix is well-draining and allows water to escape quickly. 3. Sandy soil: A sandy soil with a high percentage of sand is suitable for cactus plants. It provides excellent drainage and allows water to escape quickly. 4. Gravel and cactus mix: Layering a cactus mix on top of a layer of small rocks or gravel can also be a good option. This mimics the natural environment of cactus plants in the desert and allows water to drain away quickly. It’s essential to avoid using heavy, clay-based soils, as they can retain too much water, leading to root rot and other issues for cactus plants. Additionally, it’s essential to ensure that the soil is well-aerated, as cactus plants require good air circulation around their roots.
Online RLHF Response Cactus plants thrive in well-draining soil that allows excess water to pass through and does not hold moisture for long periods. The ideal soil for cactus plants is a mix that mimics their natural desert habitat. Here are some characteristics of suitable cactus soil: 1. **Coarse Texture**: The soil should have a coarse texture, allowing water to drain quickly. This is usually achieved by using materials like sand, perlite, or pumice. These materials help prevent waterlogged roots and root rot, which are common issues for cacti. 2. **Loamy Soil**: A loamy soil that is a combination of sand, silt, and clay is also suitable. It retains some moisture while still allowing for proper drainage. Adding some coarse materials like perlite or pumice to the loamy soil can enhance drainage further. 3. **Low Nutrient Content**: Cacti have shallow root systems and do not require a lot of nutrients. Therefore, a soil with low nutrient content, such as a cactus-specific potting mix, is ideal. Don’t use regular garden soil, as it may have too many nutrients and can lead to excessive growth and weak stems. 4. **Acidic to Slightly Alkaline**: Cactus plants prefer slightly acidic to slightly alkaline soil, with a pH range of 6.0 to 7.5. This range is generally neutral and does not require additional adjustments. 5. **Dry and Aerated**: The soil should be dry on the surface to prevent moisture buildup, but still aerated enough to allow oxygen to reach the roots. This promotes healthy root growth and prevents issues like root rot. To create a suitable cactus soil mix, you can use a combination of the following ingredients: - 2 parts cactus or succulent potting mix (which usually contains peat, perlite, and coconut coir) - 1 part coarse sand (such as sharp sand or builder’s sand) - 1 part perlite or pumice (to improve drainage) Mix these ingredients thoroughly and use the resulting soil for your cactus plants. You can also purchase commercial cactus and succulent soil mixes from a nursery or garden center, which are usually well-suited for these types of plants. Remember to always allow the soil to dry out completely between waterings, as cactus plants are adapted to survive in arid conditions. Overwatering is a common cause of cactus issues, so be cautious when watering your cacti.
Figure 12: Case Study – Example 3.